OpenStudy (anonymous):

For Module 6, solve the following system of equations by graphing, substitution,or elimination x-y=10 2x+y=2 For Module 7, simplify the following polynomials: (z+y)^2 (x-y)^2 (x+y)(x-y)

OpenStudy (amistre64):

let x=n+5 let y=n-5

OpenStudy (amistre64):

and of course "simplify" is a poor word to use inthe other one

OpenStudy (anonymous):

wht?

OpenStudy (amistre64):

i wrote: let x=n+5 let y=n-5 :/

OpenStudy (anonymous):

is that the answer

OpenStudy (amistre64):

no, the answer is the results you get after substituting that into the setup ....

OpenStudy (anonymous):

can u help me solve it and which question is this

OpenStudy (amistre64):

its the question that deals with subtitution .... or any of the other methods mentioned in it

OpenStudy (amistre64):

x-y=10 2x+y=2 let x=n+5 let y=n-5 (n+5)-(n-5)=10 2(n+5)+(n-5)=2 now solve for n

OpenStudy (anonymous):

okay so how u solve n

OpenStudy (amistre64):

solving 1 equation for 1 variable should have been gone over prior to asking you to solve for multiple equations in multiple variables .... if you have not learned how to collect like terms and such by now you may want to review past material.

OpenStudy (anonymous):

I will use the elimination method x - y = 10 2x + y = 2 ----------add 3x = 12 x = 4 now sub 4 in for x in either of the original equations 2x + y = 2 2(4) + y = 2 8 + y = 2 y = 2 - 8 y = -6 check... x - y = 10 4 - (-6) = 10 4 + 6 = 10 10 = 10 (correct) x = 4 and y = -6 ================ (z + y)^2 = (z + y)(z + y) = use the FOIL method F - multiply the first terms in each set ....z * z = z^2 O - multiply outside terms in each set ...z * y = zy I - multiply inside terms in each set ...y * z = zy L - multiply last terms in each set...y * y = y^2 now put them all together z^2 + zy + zy + y^2 z^2 + 2zy + y^2 your other problems can be solved the same way :)